a+b=20
5a+9b=120
solve for b

Answer by Boreal(15235)

Can do this by substitution or elimination. Let me substitute. a+b=20, a=20-b
5(20-b)+9b=120
100-5b+9b=120
4b=20
b=5
a=15
--------------
check in second equation
5(15)+9(5)=120, and that checks.